John can walk 4 miles in the same time it takes Paul to walk 5 miles. John, however, requires 3 mins. Longer time than Paul to walk a mile. Find their rates.


1
Expert's answer
2022-04-04T15:54:18-0400

"t_1-time\\,\\,John\\,\\,walks\\,\\,a\\,\\,mile\\\\t_2-time\\,\\,Paul\\,\\,walks\\,\\,a\\,\\,mile\\\\\\left\\{ \\begin{array}{c}\t4t_1=5t_2\\\\\tt_1=t_2+3\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\t4\\left( t_2+3 \\right) =5t_2\\\\\tt_1=t_2+3\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tt_1=15\\\\\tt_2=12\\\\\\end{array} \\right. \\\\Rates:\\\\v_1=\\frac{1}{t_1}=\\frac{1}{15}miles\/\\min \\\\v_2=\\frac{1}{t_2}=\\frac{1}{12}miles\/\\min"